OpenStudy (anonymous):

how do you find the point that the lines intersect y=-x+8 and y=4x-7

OpenStudy (anonymous):

without a graph

OpenStudy (math&ing001):

Plug in y in one of the equations, then solve for x.

OpenStudy (sleepyjess):

Without a graph? Oh, let me try to remember this

OpenStudy (math&ing001):

When you find x, plug it in any of the two equations to find y.

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

?

OpenStudy (math&ing001):

How did you get that ?

OpenStudy (sleepyjess):

You can do what @math&ing001 is saying, or you could set them equal to each other

OpenStudy (math&ing001):

-x+8 = 4x-7 We need to solve for x for this equation.

OpenStudy (anonymous):

4x-7=-x+8 5x-7=8 5x=15 x=3

OpenStudy (math&ing001):

Yeah correct !

OpenStudy (sleepyjess):

Yep, now plug in 3 for x and solve for y

OpenStudy (sleepyjess):

He can't use a graph @Nnesha

OpenStudy (anonymous):

then you just solve like a regular equation?

OpenStudy (math&ing001):

Yeah plug that value in any of the two equations to get the other coordinate of the point of intersection.

OpenStudy (sleepyjess):

@anthony81864327158756 , please don't spam others' posts with your question.

OpenStudy (anonymous):

I understand thanks

OpenStudy (sleepyjess):

Will you please delete those replies

OpenStudy (math&ing001):

welcome @Dragon3death !